OpenStudy (anonymous):

what is the equation of a line that passes through the point (0,5) and has a negative slope

OpenStudy (anonymous):

there is an infinite number of solutions actually

OpenStudy (anonymous):

you can take any slope as long as it's negative. so any equation in the form of y=ax+b where a is negative will work as long as it crosses (0,5)

OpenStudy (anonymous):

easiest solution would be y=-x+5 for example

OpenStudy (anonymous):

but y=-2x+5 would be just as good and so is y=-\(\pi\)x+5